The perimeter of a rectangle is P = 2 L + 2 W, where L is the length of the rectangle and W is the width. Find the perimeter of the rectangle if its length is 5 feet more than its width and the width is 7 feet.
Found 2 solutions by josgarithmetic, ArindamDutta1993:
Answer by josgarithmetic(39623)
Answer by ArindamDutta1993(2)

width=7 feet
length=(7+5)feet=12 feet
perimeter=2(L+W)=2(12+7) feet=38 feet.
Hence the required perimeter =38 feet